b) Suppose the magnitude of ? is 3.00 m and the angle beta = 37.0° , find the x- and y-components.
Answer #1

Let A=3.00 ,then x-component Ax=Acos(37) and y-component Ay=Asin(37)

Ax=3.00*0.79=2.37

Ay=3.00*0.60=1.80